OpenStudy (anonymous):

2^X+2^X+2^X+2^X=2^7 what is the value of the X?

OpenStudy (anonymous):

2^X+2^X+2^X+2^X = 2^7 2^(2 + X) = 2^7 (2 + X) = 7 X = 5